Libtasn1

Abstract Syntax Notation One (ASN.1) library for the GNU system

part of the GnuTLS project for version 0.3.6, 11 February 2006

Fabio Fiorina

Simon Josefsson (bug-gnutls@gnu.org)

This manual is for Libtasn1 (version 0.3.6, 11 February 2006), which is a library for Abstract Syntax Notation One (ASN.1) and Distinguish Encoding Rules (DER) manipulation. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2004, 2006 Free Software Foundation Copyright \begin{matrix}c \\ ◯ \end{matrix} 2001, 2002, 2003 Fabio Fiorina

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover Texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License.”

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

Table of Contents

1 ASN.1 structure handling . . . . . . . . . . . . . . . . . . 1

1.1 ASN.1 syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Naming. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Library Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Future developments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.1 Invoking asn1Parser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Invoking asn1Coding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Invoking asn1Decoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Function reference . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.1 ASN.1 schema functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 ASN.1 field functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 DER functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 Error handling functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.5 Auxilliary functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

Appendix A Copying This Manual . . . . . . . . . . 18

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . 18

A.1.1 ADDENDUM: How to use this License for your documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Function and Data Index . . . . . . . . . . . . . . . . . . . . . 26

1 ASN.1 structure handling

This document describes the Libtasn1 library developed for ASN.1 (Abstract Syntax Notation One) structures management.

The main features of this library are:

• On line ASN1 structure management that doesn’t require any C code file generation.

• Off line ASN1 structure management with C code file generation containing an array.

• DER (Distinguish Encoding Rules) encoding.

• No limits for INTEGER and ENUMERATED values.

• It’s Free Software. Anybody can use, modify, and redistribute the library under the terms of the GNU Lesser General Public License.

• It’s thread-safe. No global variables are used and multiple library handles and session handles may be used in parallel.

• It’s portable. It should work on all Unix like operating systems, including Windows. The library itself should be portable to any C89 system, not even POSIX is required.

1.1 ASN.1 syntax

The parser is case sensitive. The comments begin with "– " and end at the end of lines. An example is in "pkix.asn" file. ASN.1 definitions must have this syntax:

    definitions_name {<object definition>}
    DEFINITIONS <EXPLICIT or IMPLICIT> TAGS ::=
    BEGIN
    <type and constants definitions>
    END

The token "::=" must be separate from others elements, so this is a wrong declaration:

         ;; INCORRECT
         Version ::=INTEGER

the correct form is:

      Version ::= INTEGER

Here is the list of types that the parser can manage:

• INTEGER

• ENUMERATED

• BOOLEAN

• OBJECT IDENTIFIER

• NULL

• BIT STRING

• OCTET STRING • UTCTime

• GeneralizedTime

• GeneralString

• SEQUENCE

• SEQUENCE OF

• SET

• SET OF

• CHOICE

• ANY

• ANY DEFINED BY This version doesn’t manage REAL type. It doesn’t allow the "EXPORT" and "IM-

PORT" sections too.

The SIZE constraints are allowed, but no check is done on them.

1.2 Naming

Consider this definition:

    Example { 1 2 3 4 }
    DEFINITIONS EXPLICIT TAGS ::=
    BEGIN
    Group ::= SEQUENCE {
       id   OBJECT IDENTIFIER,
       value  Value
    }
    Value ::= SEQUENCE {
       value1  INTEGER,
       value2  BOOLEAN
    }
    END

To identify the type ’Group’ you have to use the null terminated string "Example.Group". These strings are used in functions that are described below.

Others examples: Field ’id’ in ’Group’ type : "Example.Group.id". Field ’value1’ in field ’value’ in type ’Group’: "Example.Group.value.value1". Elements of structured types that don’t have a name, receive the name "?1","?2", and

so on.

The name "?LAST" indicates the last element of a SET_OF or SEQUENCE_OF.

1.3 Library Notes

The header file of this library is ‘libtasn1.h’.

The main type used in it is ASN1_TYPE, and it’s used to store the ASN.1 definitions and structures (instances).

The constant ASN1 TYPE EMPTY can be used for the variable initialization. For example:

      ASN1_TYPE definitions=ASN1_TYPE_EMPTY;

Some functions require a parameter named errorDescription of char* type. The array must be already allocated and must have at least MAX_ERROR_DESCRIPTION_SIZE bytes (E.g,

as in char Description[MAX_ERROR_DESCRIPTION_SIZE];).

MAX_NAME_SIZE indicates the maximum number of characters of a name inside a file with ASN1 definitions.

1.4 Future developments

• Add functions for a C code file generation containing equivalent data structures (not a single array like now).

• Type REAL.

2 Utilities

2.1 Invoking asn1Parser

‘asn1Parser’ reads one file with ASN1 definitions and generates a file with an array to use with libtasn1 functions.

Usage:  asn1Parser [options] file
Options:
 -h : shows the help message.
 -v : shows version information and exit.
 -c : checks the syntax only.
 -o file : output file.
 -n name : array name.

2.2 Invoking asn1Coding

‘asn1Coding’ generates a DER encoding from a file with ASN1 definitions and another one with assignments.

The file with assignments must have this syntax:

InstanceName  Asn1Definition
nameString  value
nameString  value
...

The output file is a binary file with the DER encoding.

Usage:  asn1Coding [options] file1 file2
 file1 : file with ASN1 definitions.
 file2 : file with assignments.
Options:
 -h : shows the help message.
 -v : shows version information and exit.
 -c : checks the syntax only.
 -o file : output file.

2.3 Invoking asn1Decoding

‘asn1Decoding’ generates an ASN1 structure from a file with ASN1 definitions and a binary file with a DER encoding.

Usage:  asn1Decoding [options] file1 file2 type
 file1 : file with ASN1 definitions.
 file2 : binary file with a DER encoding.
 type : ASN1 definition name.
Options:
 -h : shows the help message.

-v : shows version information and exit.
-c : checks the syntax only.
-o file : output file.

3 Function reference

3.1 ASN.1 schema functions

asn1 parser2tree

[Function] asn1_retCode asn1_parser2tree (const char * file_name,
   ASN1 TYPE * definitions, char * errorDescription)

file name: specify the path and the name of file that contains ASN.1 declarations. definitions: return the pointer to the structure created from "file name" ASN.1 declarations. errorDescription: return the error description or an empty string if success. Creates the structures needed to manage the definitions included in *FILE NAME file. Returns: ASN1 SUCCESS: The file has a correct syntax and every identifier is known. ASN1 ELEMENT NOT EMPTY: *POINTER not ASN1 TYPE EMPTY. ASN1 FILE NOT FOUND: An error occured while opening FILE NAME. ASN1 SYNTAX ERROR: The syntax is not correct. ASN1 IDENTIFIER NOT FOUND: In the file there is an identifier that is not defined. ASN1 NAME TOO LONG: In the file there is an identifier whith more than MAX NAME SIZE characters.

asn1 parser2array

[Function] int asn1_parser2array (const char * inputFileName, const char *
   outputFileName, const char * vectorName, char * errorDescription)

inputFileName: specify the path and the name of file that contains ASN.1 declarations. outputFileName: specify the path and the name of file that will contain the C vector definition. vectorName: specify the name of the C vector. errorDescription: return the error description or an empty string if success. Creates a file containing a C vector to use to manage the definitions included in *INPUTFILENAME file. If *INPUTFILENAME is "/aa/bb/xx.yy" and OUTPUT-FILENAME is NULL, the file created is "/aa/bb/xx asn1 tab.c". If VECTORNAME is NULL the vector name will be "xx asn1 tab". Returns: ASN1 SUCCESS: The file has a correct syntax and every identifier is known. ASN1 FILE NOT FOUND: An error occured while opening FILE NAME. ASN1 SYNTAX ERROR: The syntax is not correct. ASN1 IDENTIFIER NOT FOUND: In the file there is an identifier that is not defined. ASN1 NAME TOO LONG: In the file there is an identifier whith more than MAX NAME SIZE characters.

3.2 ASN.1 field functions

asn1 array2tree

[Function] asn1_retCode asn1_array2tree (const ASN1 ARRAY TYPE *

   array, ASN1 TYPE * definitions, char * errorDescription)

array: specify the array that contains ASN.1 declarations definitions: return the pointer to the structure created by *ARRAY ASN.1 declarations errorDescription: return the error description. Creates the structures needed to manage the ASN.1 definitions. array is a vector

created by asn1_parser2array().

Returns: ASN1 SUCCESS: Structure created correctly. ASN1 ELEMENT NOT EMPTY: *definitions not ASN1 TYPE EMPTY. ASN1 IDENTIFIER NOT FOUND: In the file there is an identifier that is not defined (see errorDescription for more information). ASN1 ARRAY ERROR: The array pointed by array is wrong.

asn1 delete structure

[Function] asn1_retCode asn1_delete_structure (ASN1 TYPE * structure)

structure: pointer to the structure that you want to delete. Deletes the structure *structure. At the end, *structure is set to ASN1 TYPE EMPTY. Returns: ASN1 SUCCESS: Everything OK. ASN1 ELEMENT NOT FOUND: *structure was ASN1 TYPE EMPTY.

asn1 delete element

[Function] asn1_retCode asn1_delete_element (ASN1 TYPE structure, const
   char * element_name)

structure: pointer to the structure that contains the element you want to delete. element name: element’s name you want to delete. Deletes the element named *element_name inside *structure. Returns: ASN1 SUCCESS: Everything OK. ASN1 ELEMENT NOT FOUND: The name element was not found.

asn1 create element

[Function] asn1_retCode asn1_create_element (ASN1 TYPE definitions,

const char * source_name, ASN1 TYPE * element) definitions: pointer to the structure returned by "parser asn1" function source name: the name of the type of the new structure (must be inside p structure). element: pointer to the structure created. Creates a structure of type source_name. Example using "pkix.asn": rc = asn1 create structure(cert def, "PKIX1.Certificate", certptr); Returns: ASN1 SUCCESS: Creation OK. ASN1 ELEMENT NOT FOUND: SOURCE NAME isn’t known

asn1 print structure

[Function] void asn1_print_structure (FILE * out, ASN1 TYPE structure,

const char * name, int mode) out: pointer to the output file (e.g. stdout). structure: pointer to the structure that you want to visit. name: an element of the structure mode: specify how much of the structure to print, can be ASN1_PRINT_NAME, ASN1_

PRINT_NAME_TYPE, ASN1_PRINT_NAME_TYPE_VALUE, or ASN1_PRINT_ALL.

Prints on the out file descriptor the structure’s tree starting from the name element inside the structure structure.

asn1 number of elements

[Function] asn1_retCode asn1_number_of_elements (ASN1 TYPE element,

const char * name, int * num) element: pointer to the root of an ASN1 structure. name: the name of a sub-structure of ROOT. num: pointer to an integer where the result will be stored Counts the number of elements of a sub-structure called NAME with names equal to "?1","?2", ... Returns: ASN1 SUCCESS: Creation OK. ASN1 ELEMENT NOT FOUND: NAME isn’t known. ASN1 GENERIC ERROR: Pointer num equal to NULL.

asn1 find structure from oid

[Function] const char * asn1_find_structure_from_oid (ASN1 TYPE
   definitions, const char * oidValue)

definitions: ASN1 definitions oidValue: value of the OID to search (e.g. "1.2.3.4"). Search the structure that is defined just after an OID definition. Returns: NULL when OIDVALUE not found, otherwise the pointer to a constant string that contains the element name defined just after the OID.

asn1 copy node

[Function] asn1_retCode asn1_copy_node (ASN1 TYPE dst, const char *

dst_name, ASN1 TYPE src, const char * src_name) dst: Destination ASN1 TYPE node. dst name: Field name in destination node. src: Source ASN1 TYPE node. src name: Field name in source node. Create a deep copy of a ASN1 TYPE variable. Return value: Return ASN1 SUCCESS on success.

asn1 write value

[Function] asn1_retCode asn1_write_value (ASN1 TYPE node_root, const

char * name, const void * ivalue, int len) node root: pointer to a structure name: the name of the element inside the structure that you want to set. ivalue: vector used to specify the value to set. If len is >0, VALUE must be a two’s complement form integer. if len=0 *VALUE must be a null terminated string with an integer value. len: number of bytes of *value to use to set the value: value[0]..value[len-1] or 0 if value is a null terminated string Set the value of one element inside a structure. If an element is OPTIONAL and you want to delete it, you must use the value=NULL and len=0. Using "pkix.asn": result=asn1 write value(cert, "tbsCertificate.issuerUniqueID", NULL, 0); Description for each type: INTEGER: VALUE must contain a two’s complement form integer. value[0]=0xFF , len=1 -> integer=-1. value[0]=0xFF value[1]=0xFF , len=2 -> integer=-1. value[0]=0x01 , len=1 -> integer= 1. value[0]=0x00 value[1]=0x01 , len=2 -> integer= 1. value="123" , len=0 -> integer= 123. ENUMERATED: As INTEGER (but only with not negative numbers). BOOLEAN: VALUE must be the null terminated string "TRUE" or "FALSE" and LEN != 0. value="TRUE" , len=1 -> boolean=TRUE. value="FALSE" , len=1 -> boolean=FALSE. OBJECT IDENTIFIER: VALUE must be a null terminated string with each number separated by a dot (e.g. "1.2.3.543.1"). LEN != 0. value="1 2 840 10040 4 3" , len=1 -> OID=dsa-with-sha. UTCTime: VALUE must be a null terminated string in one of these formats: "YYMMDDhhmmssZ", "YYMMDDhhmmssZ", "YYMMDDhhmmss+hh’mm’", "YYMMDDhhmmss-hh’mm’", "YYMMDDhhmm+hh’mm’", or "YYMMDDhhmmhh’mm’". LEN != 0. value="9801011200Z" , len=1 -> time=Jannuary 1st, 1998 at 12h 00m Greenwich Mean Time GeneralizedTime: VALUE must be in one of this format: "YYYYMMDDhh-

mmss.sZ",  "YYYYMMDDhhmmss.sZ",  "YYYYMMDDhhmmss.s+hh’mm’",
"YYYYMMDDhhmmss.s-hh’mm’",  "YYYYMMDDhhmm+hh’mm’",  or

"YYYYMMDDhhmm-hh’mm’" where ss.s indicates the seconds with any precision like "10.1" or "01.02". LEN != 0 value="2001010112001.12-0700" , len=1 -> time=Jannuary 1st, 2001 at 12h 00m 01.12s Pacific Daylight Time OCTET STRING: VALUE contains the octet string and LEN is the number of octets. value="$\backslash$x01$\backslash$x02$\backslash$x03" , len=3 -> three bytes octet string GeneralString: VALUE contains the generalstring and LEN is the number of octets. value="$\backslash$x01$\backslash$x02$\backslash$x03" , len=3 -> three bytes generalstring BIT STRING: VALUE contains the bit string organized by bytes and LEN is the number of bits. value="$\backslash$xCF" , len=6 -> bit string="110011" (six bits) CHOICE: if NAME indicates a choice type, VALUE must specify one of the alternatives with a null terminated string. LEN != 0. Using "pkix.asn"\: result=asn1 write value(cert, "certificate1.tbsCertificate.subject", "rdnSequence", 1); ANY: VALUE indicates the der encoding of a structure. LEN != 0. SEQUENCE OF: VALUE must be the null terminated string "NEW" and LEN != 0. With this instruction another element is appended in the sequence. The name of this element will be "?1" if it’s the first one, "?2" for the second and so on. Using "pkix.asn"\: result=asn1 write value(cert, "certificate1.tbsCertificate.subject.rdnSequence", "NEW", 1); SET OF: the same as SEQUENCE OF. Using "pkix.asn": result=asn1 write value(cert, "tbsCertificate.subject.rdnSequence.?LAST", "NEW", 1); Returns: ASN1 SUCCESS: Set value OK. ASN1 ELEMENT NOT FOUND: NAME is not a valid element. ASN1 VALUE NOT VALID: VALUE has a wrong format.

asn1 read value

[Function] asn1_retCode asn1_read_value (ASN1 TYPE root, const char *
   name, void * ivalue, int * len)

root: pointer to a structure. name: the name of the element inside a structure that you want to read. ivalue: vector that will contain the element’s content, must be a pointer to memory cells already allocated. len: number of bytes of *value: value[0]..value[len-1]. Initialy holds the sizeof value. Returns the value of one element inside a structure.

If an element is OPTIONAL and the function "read value" returns ASN1_ELEMENT_ NOT_FOUND, it means that this element wasn’t present in the der encoding that created the structure. The first element of a SEQUENCE OF or SET OF is named "?1". The second one "?2" and so on. INTEGER: VALUE will contain a two’s complement form integer. integer=-1 -> value[0]=0xFF , len=1. integer=1 -> value[0]=0x01 , len=1. ENUMERATED: As INTEGER (but only with not negative numbers). BOOLEAN: VALUE will be the null terminated string "TRUE" or "FALSE" and LEN=5 or LEN=6. OBJECT IDENTIFIER: VALUE will be a null terminated string with each number separated by a dot (i.e. "1.2.3.543.1"). LEN = strlen(VALUE)+1 UTCTime: VALUE will be a null terminated string in one of these formats: "YYM-MDDhhmmss+hh’mm’" or "YYMMDDhhmmss-hh’mm’". LEN=strlen(VALUE)+1. GeneralizedTime: VALUE will be a null terminated string in the same format used to set the value. OCTET STRING: VALUE will contain the octet string and LEN will be the number of octets. GeneralString: VALUE will contain the generalstring and LEN will be the number of octets. BIT STRING: VALUE will contain the bit string organized by bytes and LEN will be the number of bits. CHOICE: If NAME indicates a choice type, VALUE will specify the alternative selected. ANY: If NAME indicates an any type, VALUE will indicate the DER encoding of the structure actually used. Returns: ASN1 SUCCESS: Set value OK. ASN1 ELEMENT NOT FOUND: NAME is not a valid element. ASN1 VALUE NOT FOUND: There isn’t any value for the element selected. ASN1 MEM ERROR: The value vector isn’t big enough to store the result. In this case LEN will contain the number of bytes needed.

asn1 read tag

[Function] asn1_retCode asn1_read_tag (node asn * root, const char * name, int

   * tagValue, int * classValue)

root: pointer to a structure name: the name of the element inside a structure. tagValue: variable that will contain the TAG value. classValue: variable that will specify the TAG type. Returns the TAG and the CLASS of one element inside a structure.

CLASS can have one of these constants: ASN1_CLASS_APPLICATION, ASN1_CLASS_
UNIVERSAL, ASN1_CLASS_PRIVATE or ASN1_CLASS_CONTEXT_SPECIFIC.

Returns: ASN1 SUCCESS: Set value OK. ASN1 ELEMENT NOT FOUND: NAME is not a valid element.

3.3 DER functions

asn1 length der

[Function] void asn1_length_der (unsigned long int len, unsigned char * ans, int

   * ans_len)

len: value to convert. ans: string returned. ans len: number of meaningful bytes of ANS (ans[0]..ans[ans len-1]). Creates the DER coding for the LEN parameter (only the length). The ans buffer is pre-allocated and must have room for the output.

asn1 octet der

[Function] void asn1_octet_der (const unsigned char * str, int str_len,

unsigned char * der, int * der_len) str: OCTET string. str len: STR length (str[0]..str[str len-1]). der: string returned. der len: number of meaningful bytes of DER (der[0]..der[ans len-1]). Creates the DER coding for an OCTET type (length included).

asn1 bit der

[Function] void asn1_bit_der (const unsigned char * str, int bit_len, unsigned

char * der, int * der_len) str: BIT string. bit len: number of meaningful bits in STR. der: string returned. der len: number of meaningful bytes of DER (der[0]..der[ans len-1]). Creates the DER coding for a BIT STRING type (length and pad included).

asn1 der coding

[Function] asn1_retCode asn1_der_coding (ASN1 TYPE element, const char *
   name, void * ider, int * len, char * ErrorDescription)

element: pointer to an ASN1 element name: the name of the structure you want to encode (it must be inside *POINTER). ider: vector that will contain the DER encoding. DER must be a pointer to memory cells already allocated. len: number of bytes of *ider: ider[0]..ider[len-1], Initialy holds the sizeof of der vector.

Creates the DER encoding for the NAME structure (inside *POINTER structure). Returns: ASN1 SUCCESS: DER encoding OK. ASN1 ELEMENT NOT FOUND: NAME is not a valid element. ASN1 VALUE NOT FOUND: There is an element without a value. ASN1 MEM ERROR: ider vector isn’t big enough. Also in this case LEN will contain the length needed.

asn1 get length der

[Function] signed long asn1_get_length_der (const unsigned char * der, int

   der_len, int * len)

der: DER data to decode. der len: Length of DER data to decode. len: Output variable containing the length of the DER length field. Extract a length field from DER data. Return value: Return the decoded length value, or -1 on indefinite length, or -2 when the value was too big.

asn1 get tag der

[Function] int asn1_get_tag_der (const unsigned char * der, int der_len,

unsigned char * class, int * len, unsigned long * tag) der: DER data to decode. der len: Length of DER data to decode. class: Output variable containing decoded class. len: Output variable containing the length of the DER TAG data. tag: Output variable containing the decoded tag. Decode the class and TAG from DER code. Return value: Returns ASN1 SUCCESS on success, or an error.

asn1 get octet der

[Function] int asn1_get_octet_der (const unsigned char * der, int der_len, int

   * ret_len, unsigned char * str, int str_size, int * str_len)

der: DER data to decode containing the OCTET SEQUENCE. der len: Length of DER data to decode. ret len: Output variable containing the length of the DER data. str: Pre-allocated output buffer to put decoded OCTET SEQUENCE in. str size: Length of pre-allocated output buffer. str len: Output variable containing the length of the OCTET SEQUENCE. Extract an OCTET SEQUENCE from DER data. Return value: Returns ASN1 SUCCESS on success, or an error.

asn1 get bit der

[Function] int asn1_get_bit_der (const unsigned char * der, int der_len, int *

   ret_len, unsigned char * str, int str_size, int * bit_len)

der: DER data to decode containing the BIT SEQUENCE. der len: Length of DER data to decode. ret len: Output variable containing the length of the DER data. str: Pre-allocated output buffer to put decoded BIT SEQUENCE in. str size: Length of pre-allocated output buffer. bit len: Output variable containing the size of the BIT SEQUENCE. Extract a BIT SEQUENCE from DER data. Return value: Return ASN1 SUCCESS on success, or an error.

asn1 der decoding

[Function] asn1_retCode asn1_der_decoding (ASN1 TYPE * element, const
   void * ider, int len, char * errorDescription)

element: pointer to an ASN1 structure. ider: vector that contains the DER encoding. len: number of bytes of *ider: ider[0]..ider[len-1]. errorDescription: null-terminated string contains details when an error occurred. Fill the structure *ELEMENT with values of a DER encoding string. The sructure must just be created with function ’create stucture’. If an error occurs during the decoding procedure, the *ELEMENT is deleted and set equal to ASN1_TYPE_EMPTY. Returns: ASN1 SUCCESS: DER encoding OK. ASN1 ELEMENT NOT FOUND: ELEMENT is ASN1 TYPE EMPTY. ASN1 TAG ERROR,ASN1 DER ERROR: The der encoding doesn’t match the structure NAME. *ELEMENT deleted.

asn1 der decoding element

[Function] asn1_retCode asn1_der_decoding_element (ASN1 TYPE *

structure, const char * elementName, const void * ider, int len, char *

   errorDescription)

structure: pointer to an ASN1 structure elementName: name of the element to fill ider: vector that contains the DER encoding of the whole structure. len: number of bytes of *der: der[0]..der[len-1] errorDescription: null-terminated string contains details when an error occurred. Fill the element named ELEMENTNAME with values of a DER encoding string. The sructure must just be created with function ’create stucture’. The DER vector must contain the encoding string of the whole STRUCTURE. If an error occurs during the decoding procedure, the *STRUCTURE is deleted and set equal to ASN1_TYPE_

EMPTY.

Returns: ASN1 SUCCESS: DER encoding OK. ASN1 ELEMENT NOT FOUND: ELEMENT is ASN1 TYPE EMPTY or element-Name == NULL. ASN1 TAG ERROR,ASN1 DER ERROR: The der encoding doesn’t match the structure STRUCTURE. *ELEMENT deleted.

asn1 der decoding startEnd

[Function] asn1_retCode asn1_der_decoding_startEnd (ASN1 TYPE

element, const void * ider, int len, const char * name_element, int *

   start, int * end)

element: pointer to an ASN1 element ider: vector that contains the DER encoding. len: number of bytes of *ider: ider[0]..ider[len-1] name element: an element of NAME structure. start: the position of the first byte of NAME ELEMENT decoding (ider[*start]) end: the position of the last byte of NAME ELEMENT decoding (ider[*end]) Find the start and end point of an element in a DER encoding string. I mean that if you have a der encoding and you have already used the function "asn1 der decoding" to fill a structure, it may happen that you want to find the piece of string concerning an element of the structure. Example: the sequence "tbsCertificate" inside an X509 certificate. Returns: ASN1 SUCCESS: DER encoding OK. ASN1 ELEMENT NOT FOUND: ELEMENT is ASN1 TYPE EMPTY or NAME ELEMENT is not a valid element. ASN1 TAG ERROR,ASN1 DER ERROR: the der encoding doesn’t match the structure ELEMENT.

asn1 expand any defined by

[Function] asn1_retCode asn1_expand_any_defined_by (ASN1 TYPE
   definitions, ASN1 TYPE * element)

definitions: ASN1 definitions element: pointer to an ASN1 structure Expands every "ANY DEFINED BY" element of a structure created from a DER decoding process (asn1 der decoding function). The element ANY must be defined by an OBJECT IDENTIFIER. The type used to expand the element ANY is the first one following the definition of the actual value of the OBJECT IDENTIFIER. Returns: ASN1 SUCCESS: Substitution OK. ASN1 ERROR TYPE ANY: Some "ANY DEFINED BY" element couldn’t be expanded due to a problem in OBJECT ID -> TYPE association. other errors: Result of der decoding process.

asn1 expand octet string

[Function] asn1_retCode asn1_expand_octet_string (ASN1 TYPE

definitions, ASN1 TYPE * element, const char * octetName, const char

   * objectName)

definitions: ASN1 definitions element: pointer to an ASN1 structure octetName: name of the OCTECT STRING field to expand. objectName: name of the OBJECT IDENTIFIER field to use to define the type for expansion. Expands an "OCTET STRING" element of a structure created from a DER decoding process (asn1 der decoding function). The type used for expansion is the first one following the definition of the actual value of the OBJECT IDENTIFIER indicated by OBJECTNAME. Returns: ASN1 SUCCESS: Substitution OK. ASN1 ELEMENT NOT FOUND: OBJECTNAME or OCTETNAME are not correct. ASN1 VALUE NOT VALID: Wasn’t possible to find the type to use for expansion. other errors: result of der decoding process.

3.4 Error handling functions

libtasn1 perror

[Function] void libtasn1_perror (asn1 retCode error)

error: is an error returned by a libtasn1 function. This function is like perror(). The only difference is that it accepts an error returned by a libtasn1 function.

libtasn1 strerror

[Function] const char * libtasn1_strerror (asn1 retCode error)

error: is an error returned by a libtasn1 function. This function is similar to strerror(). The only difference is that it accepts an error (number) returned by a libtasn1 function. Returns: Pointer to static zero-terminated string describing error code.

3.5 Auxilliary functions

asn1 find node

[Function] ASN1_TYPE asn1_find_node (ASN1 TYPE pointer, const char *
   name)

pointer: NODE ASN element pointer. name: null terminated string with the element’s name to find.

Searches for an element called NAME starting from POINTER. The name is composed by differents identifiers separated by dots. When *POINTER has a name, the first identifier must be the name of *POINTER, otherwise it must be the name of one child of *POINTER. Return value: the searching result. NULL if not found.

asn1 check version

[Function] const char * asn1_check_version (const char * req_version)

req version: Required version number, or NULL. Check that the the version of the library is at minimum the requested one and return the version string; return NULL if the condition is not satisfied. If a NULL is passed to this function, no check is done, but the version string is simply returned.

See LIBTASN1_VERSION for a suitable req_version string.

Return value: Version string of run-time library, or NULL if the run-time library does not meet the required version number.

Appendix A Copying This Manual

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

A.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 25

Concept Index

A

ASN.1 schema . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 asn1Coding program . . . . . . . . . . . . . . . . . . . . . . . . . . 4 asn1Decoding program . . . . . . . . . . . . . . . . . . . . . . . . 4 asn1Parser program . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

F

FDL, GNU Free Documentation License . . . . . . . 18 Future developments . . . . . . . . . . . . . . . . . . . . . . . . . . 3

H

Header file libtasn1.h. . . . . . . . . . . . . . . . . . . . . . . . . . 3

M

Main type ASN1 TYPE . . . . . . . . . . . . . . . . . . . . . . . 3

P

Porting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

S

Supported ASN.1 types, list of . . . . . . . . . . . . . . . . . 1

T

threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Function and Data Index 26

Function and Data Index

A

asn1_array2tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 asn1_bit_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 asn1_check_version . . . . . . . . . . . . . . . . . . . . . . . . . 17 asn1_copy_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_create_element . . . . . . . . . . . . . . . . . . . . . . . . . 7 asn1_delete_element . . . . . . . . . . . . . . . . . . . . . . . . . 7 asn1_delete_structure. . . . . . . . . . . . . . . . . . . . . . . 7 asn1_der_coding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 asn1_der_decoding . . . . . . . . . . . . . . . . . . . . . . . . . . 14

asn1_der_decoding_element . . . . . . . . . . . . . . . . . 14
asn1_der_decoding_startEnd . . . . . . . . . . . . . . . . 15
asn1_expand_any_defined_by . . . . . . . . . . . . . . . . 15
asn1_expand_octet_string . . . . . . . . . . . . . . . . . . 16

asn1_find_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

asn1_find_structure_from_oid . . . . . . . . . . . . . . . 8

asn1_get_bit_der . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 asn1_get_length_der. . . . . . . . . . . . . . . . . . . . . . . . 13 asn1_get_octet_der . . . . . . . . . . . . . . . . . . . . . . . . . 13 asn1_get_tag_der . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 asn1_length_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

                                                asn1_number_of_elements . . . . . . . . . . . . . . . . . . . . 8

asn1_octet_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 asn1_parser2array . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 asn1_parser2tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 asn1_print_structure. . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_read_tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 asn1_read_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 asn1_write_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

L

libtasn1_perror . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 libtasn1_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . 16